OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

@tafkas77

OpenStudy (anonymous):

Okay. This one is going to be a LOT like the one we just covered. First things first, though: Do you know what parallel lines have in common?

OpenStudy (anonymous):

Yeah I think so.

OpenStudy (anonymous):

Well, just to be sure... :) Parallel lines have the SAME slope. So if we're looking for an equation that has a line parallel to the equation they gave you, all we need to have is the same slope! In this case, what would our slope be? (Hint: Look for the "m" in y = 7x + 2)

OpenStudy (anonymous):

the m is 7

OpenStudy (anonymous):

Perfect. So now we can move on to our list of steps. What should we do first, since we know our slope and our equation (y = mx + b)?

OpenStudy (anonymous):

we plug in the slope.

OpenStudy (anonymous):

Right. :) y = 7x + b Okay, what should our next step look like?

OpenStudy (anonymous):

we plug in the x and y intercepts. -3 = 71 + b?

OpenStudy (anonymous):

oops -3 +7(1) +b

OpenStudy (anonymous):

Correct. :) Now, we have to simplify the equation: -3 = 7 + b What happens next?

OpenStudy (anonymous):

-3 = 7 + b

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

we get b alone.

OpenStudy (anonymous):

Exactly. And how do you do that?

OpenStudy (anonymous):

I dont know. im a little stuck.

OpenStudy (anonymous):

That's all right. :) Well, to get 'b' alone, you want to CANCEL OUT any numbers that 'b' is with. :) Let's look at what we've got... -3 = 7 + b Well, we need to get rid of 7. We want to banish 7 to the other side of the equation. But how?

OpenStudy (anonymous):

i dont know.

OpenStudy (anonymous):

We'll add a -7 to both sides, because whatever you do to one side, you do to the other. -7 - 3 = 7 - 7 + b

OpenStudy (anonymous):

7 plus -7 equals zero, so that means 'b' gets a little fresh air, know what I mean? Haha, does this make sense?

OpenStudy (anonymous):

yeah it does.

OpenStudy (anonymous):

Great! So can you simplify -7 - 3 = 7 - 7 + b

OpenStudy (anonymous):

-10 = 0 + b

OpenStudy (anonymous):

Perfect. So b = -10. Now we know both the slope (7) and the y-intercept (b). Can you solve the original equation (y = mx + b) now?

OpenStudy (anonymous):

so the answer is y= 7 + -10

OpenStudy (anonymous):

y = 7 -10

OpenStudy (anonymous):

Very close. Hey, whenever you see y = mx + b think of it as y = mx +/- b, okay? :) And leave the x in there, so you're still following the format (like you did last time): y = 7x - 10 Awesome work!

OpenStudy (anonymous):

thanks for your help! :D

OpenStudy (anonymous):

No problem! (Pun totally intended, hehe) :)